# Lindblad quantum dynamics from correlation functions of classical spin chains

### v1

@@ Abstract
The Lindblad quantum master equation is one of the central approaches to the physics of open quantum systems. In particular, boundary driving enables the study of transport, where a steady state emerges in the long-time limit, which features a constant current and a characteristic density profile. While the Lindblad equation complements other approaches to transport in closed quantum systems, it has become clear that a connection between closed and open systems exists in certain cases. Here, we build on this connection for magnetization transport in the spin-1/2 XXZ chain with and without integrability-breaking perturbations. Specifically, we show that the time evolution of the open quantum system can be described on the basis of classical correlation functions, as generated by the Hamiltonian equations of motion for real vectors. By comparing to exact numerical simulations of the Lindblad equation, we demonstrate the accuracy of this description for a range of model parameters, but also point out counterexamples. While this agreement is an interesting physical observation, it also suggests that classical mechanics can be used to solve the Lindblad equation for comparatively large system sizes, which lie outside the possibilities of a quantum mechanical treatment.
@@ I Introduction
Within the diverse class of nonequilibrium processes, transport is a natural one for systems with one or more globally conserved quantities [21], like total energy, particle number, or magnetization. Transport is further a process which is relevant to both, closed and open situations. In an open situation, the system of actual interest can be coupled at its boundaries to two reservoirs at different temperatures or chemical potentials, such that transport is induced and a nonequilibrium steady state is usually established in the long-time limit [22, 23, 24]. In this steady state, the constant current and the form of the density profile yield information on the qualitative type of transport and also allow to determine quantitative values for transport coefficients [21]. A widely used strategy to describe such an open scenario is the Lindblad equation [25], which has its assets and drawbacks. On the one hand, the derivation of this equation from a microscopic system-bath model can be a nontrivial task in praxis [26, 27]. On the other hand, it is the most general form of a quantum master equation, which is local in time and maps any density matrix to a density matrix again. Furthermore, the structure of the Lindblad equation enables the application of well-suited numerical techniques. Here, one method is provided by the concept of stochastic unraveling [28, 29], which constructs the time evolution of the density matrix in Liouville space as the average over many pure-state trajectories in Hilbert space. An alternative method is given by matrix product states [30, 31, 23, 32], where entanglement growth is reduced because of dissipation.
@@
In this paper, we also investigate to what extent the time evolution in a quantum system can be described on the basis of the dynamics in the classical counterpart. In contrast to previous works, which have been devoted to a comparison of the corresponding correlation functions in a closed scenario [64, 65], we intend to go a substantial step beyond. Specifically, we are going to compare the open quantum system to the closed classical system. For such a comparison, we obviously need a connection between the dynamics in open and closed scenarios [67, 68], which does not exist in general [69, 70, 71]. For the spin-$1/2$ chain, however, a connection has been recently suggested [72] for the case of small system-bath coupling and weak driving. Because this connection involves quantum correlation functions, we replace them by the corresponding classical ones. In this way, we can address the main question of our work: Is it possible to obtain Lindblad quantum dynamics from correlation functions of classical spin chains? To answer this question, we compare to exact numerical simulations of the Lindblad equation. While we observe a remarkably good agreement for a range of model parameters, we also point out counterexamples. This agreement further hints that classical mechanics can be used as a strategy to solve the Lindblad equation for large system sizes, which are not accessible in a quantum mechanical treatment. Such a classical strategy has been also discussed in other open quantum systems [73].
@@ II Open quantum system
as the most general form of a quantum master equation, which is local in time and maps any density matrix to a density matrix again [25]. While the first term on the r.h.s. is coherent and describes the unitary time evolution w.r.t. to a given Hamiltonian $H$, the second term on the r.h.s. is incoherent and describes the damping due to the presence of a bath. This damping reads
@@
An exact analytical solution of the Lindblad equation, or an accurate approximation of it, can in general not be derived. Hence, one has to resort typically to numerical methods. In this context, standard exact diagonalization is particularly challenging, since the Liouville space (of dimension $D=2^{N}\times 2^{N})$ is substantially larger than the anyhow large Hilbert space ($D=2^{N}$). Yet, the Lindblad form allows for stochastic unraveling [28, 29], which yields the time-dependent density matrix as the average over many pure-state trajectories. Additionally, simulations on the basis of matrix product states [30, 31, 23, 32] give access to systems of hundreds of spins, at least on time scales with a still low amount of entanglement.
@@ III Connection between open and closed systems
where the sum runs over $T_{\text{max}}$ different time sequences $(\tau_{1},\tau_{2},\ldots)$. Here, a particular time sequence is generated by
@@
where $\varepsilon_{j+1}$ are random numbers drawn from a uniform distribution in the interval ]0,1]]0,1]] 0 , 1 ]. The amplitudes $A_{j}$ are now given by
@@
If $d_{B_{1}}(\tau_{j}-0^{+})\to 0$, $A_{j}\to 1/2$.
While the connection might look quite complicated at first sight, it is rather simple, especially since it expresses the nonequilibrium dynamics in the open system just as a superposition of equilibrium correlation functions in the closed system. As discussed in Ref. [72], such a connection cannot always hold and requires sufficiently small values of both, $\gamma$ and $\mu$. Furthermore, it assumes the absence of nondecaying edge modes, which can occur in a closed system with open boundary conditions [75, 76]. For this reason, we focus on periodic boundary conditions.
@@
In Fig. 2, we illustrate the quality of the connection, by comparing the prediction on the r.h.s. of Eq. (15) to a simulation based on stochastic unraveling for the l.h.s. of Eq. (15). We do so for the model $H^{\prime}$ in Eq. (4) with $\Delta=1.5$, $\Delta^{\prime}=0$, $N=20$, periodic boundary conditions, small coupling $\gamma/J=0.1$, and weak driving $\mu=0.1$. For this set of parameters, the agreement between both sides is remarkable. While Fig. 2 shows existing data from the literature [72], it already depicts a prediction by the use of classical mechanics, as a main result of our work. This prediction is discussed in the following.
@@ IV Classical limit
with the Kronecker symbol $\delta_{rr^{\prime}}$ and the antisymmetric Levi-Civita tensor $\epsilon_{ijk}$. Here, we write $\hbar$ explicitly, while it is otherwise set to unity. The classical counterpart of our models now results by taking the limit
$S\to\infty\,,\quad\hbar\to 0\,,\quad\hbar\sqrt{S(S+1)}=\text{const.}$. (20)
@@
$\frac{\text{d}}{\text{d}t}\mathbf{S}_{r}=\frac{\partial H}{\partial\mathbf{S}_{r}}\times\mathbf{S}_{r}\,.$ (22)
Physically, these equations describe the precession of a spin around a magnetic field, which is generated by the interaction with the neighboring spins. Mathematically, they are a coupled set of nonlinear differential equations, which is nonintegrable by means of the Liouville-Arnold theorem, even for $\Delta^{\prime}=0$ [77, 78]. Therefore, for nontrivial initial conditions, they have to be solved numerically, as we also do here. Throughout our work, we employ a fourth-order Runge-Kutta scheme with a time step $\delta tJ=0.01$, which is small enough to ensure that the total magnetization is well conserved during the time evolution.
@@
where $R_{\text{max}}$ is the number of realizations for different initial conditions, which are randomly drawn from phase space. Formally, $R_{\text{max}}\to\infty$. In praxis, we average over as many realizations as $R_{\text{max}}={\cal O}(10^{8}-10^{9})$, to ensure that the remaining stochastic fluctuations are low. The need for such an extensive averaging in the numerical simulations is kind of compensated by the fact that the phase space grows only linearly with $N$, in contrast to the exponential growth of the Hilbert space. Due to this fact, we can particularly treat classical systems of quite large size $N$, which lie outside the possibilities of a quantum mechanical treatment. Still, we also consider small $N$ to enable a comparison to quantum mechanics for the same size.
@@
Eventually, coming back to the connection between the dynamics in open and closed systems, Eq. (15), the main idea is to replace the quantum correlation functions by the corresponding classical ones. When we do this strong simplification, the key question of our work is whether or not the time evolution in the open system can be still described accurately. To answer this question in a kind of fair way, we first need to rescale the classical time by a factor
@@ V.1 Nearest-neighbor interactions
Using the correlations in Fig. 3, we now calculate the corresponding predictions for the time evolution of the open system, according to Eq. (15). In this calculation, we average over $T_{\text{max}}={\cal O}(10^{4})$ different time sequences, which is enough to obtain sufficiently smooth curves. As already advertised before, Fig. 2 illustrates for $\Delta=1.5$ a convincing agreement with existing simulations [72] on the basis of stochastic unraveling. In Fig. 4, we also depict a comparison for $\Delta=1.0$ and $0.5$. While the agreement for $\Delta=1.0$ is also good, some disagreement is visble for $\Delta=0.5$. Of course, one might expect this disagreement in view of the differences between quantum and classical correlations in Fig. 3 (third row). However, it should be noted that also the quantum prediction differs from the stochastic-unravling data. This deviation can be traced back to the fact that the ballistic quantum motion for $\Delta=0.5$ partially violates the equilibration assumption, which underlies the derivation of Eq. (15).
@@ V.3 Larger system sizes
In Fig. 7 (c), we additionally depict a quantity, which we have not discussed so far. This quantity is the induced magnetization by the first bath, $\langle\delta S^{z}_{B_{1}}(t)\rangle$, as defined in the appendix. The induced magnetization is of interest, since from its time derivative, d/$\text{d}/\text{d}t\,\langle\delta S^{z}_{B_{1}}(t)\rangle$, the current in the steady state and then a quantitative value for the diffusion constant can be calculated. As can be seen in Fig. 7 (c), the predictions for $\langle\delta S^{z}_{B_{1}}(t)\rangle$ are kind of close to each other. Yet, they are clearly different. Hence, we obtain two different diffusion coefficients,
$D_{Q}/J\approx 0.99\,,\quad D_{C}/J\approx 0.53\,.$ (25)
Figure 9: Similar data as the one in Fig. 7, but now for a large system size $N=200$. For this $N$, only classical mechanics is feasible.
Because the prediction for the injected magnetization seems to be more sensitive to short-time deviations, as the ones visible in Fig. 3, we redo the analysis in Fig. 7 for $\gamma/J=0.05$ (instead of $\gamma/J=0.1$). In this way, the relevant time scale is increased. Simultaneously, we use another $\mu=0.05$ (instead of $\mu=0.1$). And indeed, as shown in Fig. 8 (c), the predictions for $\langle\delta S^{z}_{B_{1}}(t)\rangle$ become closer to each other, which also leads to a slightly better agreement of the diffusion constants,
$D_{Q}/J\approx 1.01\,,\quad D_{C}/J\approx 0.61\,.$ (26)
This observation indicates that one has to devote special attention to the role of $\gamma$ and $\mu$, to extract quantitative values for transport constants.
It is instructive to compare the values for the diffusion constants in Eq. (27) to other values in the literature. In the classical case, a value of $D_{C}/J\sim 0.6$ has been found for the closed system [47], which agrees nicely with the value for the open system in Eq. (27). In the quantum case, the value is known with less precision and might be $D_{Q}/J\sim 0.6$ or larger in the closed system [21], while a value of $D_{Q}/J\approx 0.58$ has been found in the open system [23, 24], based on matrix-product-state simulations of the Lindblad equation, yet for a larger $\gamma/J=1$.
@@
$D_{C}/J\approx 0.78\,,$ (27)
which is comparable to but larger than $D_{C}/J\approx 0.53$ for $N=36$, cf. Eq. (25), and indicates the importance of a finite-size analysis for the quantitative extraction of transport coefficients.

### v2

@@ Abstract
The Lindblad quantum master equation is one of the central approaches to the physics of open quantum systems. In particular, boundary driving enables the study of transport, where a steady state emerges in the long-time limit, which features a constant current and a characteristic density profile. While the Lindblad equation complements other approaches to transport in closed quantum systems, it has become clear that a connection between closed and open systems exists in certain cases. Here, we build on this connection for magnetization transport in the spin-1/2 XXZ chain with and without integrability-breaking perturbations. Specifically, we study the question whether the time evolution of the open quantum system can be described on the basis of classical correlation functions, as generated by the Hamiltonian equations of motion for real vectors. By comparing to exact numerical simulations of the Lindblad equation, we find a good accuracy of such a description for a range of model parameters, which is consistent with previous studies on closed systems. While this agreement is an interesting physical observation, it also suggests that classical mechanics can be used to solve the Lindblad equation for comparatively large system sizes, which lie outside the possibilities of a quantum mechanical treatment. We also point out counterexamples and limitations for the quantitative extraction of transport coefficients. Remarkably, our classical approach to large open systems allows to detect superdiffusion at the isotropic point.
@@ I Introduction
Within the diverse class of nonequilibrium processes, transport is a natural one for systems with one or more globally conserved quantities [21], like total energy, particle number, or magnetization. Transport is further a process which is relevant to both, closed and open situations. In an open situation, the system of actual interest can be coupled at its boundaries to two reservoirs at different temperatures or chemical potentials, such that transport is induced and a nonequilibrium steady state is usually established in the long-time limit [22, 23, 24]. In this steady state, the constant current and the form of the density profile yield information on the qualitative type of transport and also allow to determine quantitative values for transport coefficients [21]. A widely used strategy to describe such an open scenario is the Lindblad equation [25], which has its assets and drawbacks. On the one hand, the derivation of this equation from a microscopic system-bath model can be a nontrivial task in praxis [26, 27]. On the other hand, it is the most general form of a quantum master equation, which Markovian and maps any density matrix to a density matrix again. Furthermore, the structure of the Lindblad equation enables the application of well-suited numerical techniques. Here, one method is provided by the concept of stochastic unraveling [28, 29], which constructs the time evolution of the density matrix in Liouville space as the average over many pure-state trajectories in Hilbert space. An alternative method is given by matrix product states [30, 31, 23, 32], where entanglement growth is reduced because of dissipation.
@@
In this paper, we also investigate to what extent the time evolution in a quantum system can be described on the basis of the dynamics in the classical counterpart. In contrast to previous works, which have been devoted to a comparison of the corresponding correlation functions in a closed scenario [64, 65], we intend to go a substantial step beyond. Specifically, we are going to compare the open quantum system to the closed classical system. For such a comparison, we obviously need a connection between the dynamics in open and closed scenarios [67, 68], which does not exist in general [69, 70, 71]. For the spin-$1/2$ chain, however, a connection has been recently suggested [72] for the case of small system-bath coupling and weak driving. Because this connection involves quantum correlation functions, we replace them by the corresponding classical ones. In this way, we can address the main question of our work: Is it possible to obtain Lindblad quantum dynamics from correlation functions of classical spin chains? To answer this question, we compare to exact numerical simulations of the Lindblad equation. We observe a good agreement for a range of model parameters, which is consistent with previous studies on closed systems. This agreement further hints that classical mechanics can be used as a strategy to solve the Lindblad equation for large system sizes, which are not accessible in a quantum mechanical treatment. Such a classical strategy has been also discussed in other open quantum systems [73]. We also point out counterexamples and limitations for the quantitative extraction of transport coefficients. Remarkably, our classical approach to large open systems allows to detect superdiffusion at the isotropic point.
@@ II Open quantum system
as the most general form of a quantum master equation, which is Markovian and maps any density matrix to a density matrix again [25]. While the first term on the r.h.s. is coherent and describes the unitary time evolution w.r.t. to a given Hamiltonian $H$, the second term on the r.h.s. is incoherent and describes the damping due to the presence of a bath. This damping reads
@@
An exact analytical solution of the Lindblad equation, or an accurate approximation of it, can in general not be derived. Hence, one has to resort typically to numerical methods. In this context, standard exact diagonalization is particularly challenging [74], since the Liouville space (of dimension $D=2^{N}\times 2^{N})$ is substantially larger than the anyhow large Hilbert space ($D=2^{N}$). Yet, the Lindblad form allows for stochastic unraveling [28, 29], which yields the time-dependent density matrix as the average over many pure-state trajectories. Additionally, simulations on the basis of matrix product states [30, 31, 23, 32] give access to systems of hundreds of spins, at least on time scales with a still low amount of entanglement.
@@ III Connection between open and closed systems
where the sum runs over $T_{\text{max}}$ different time sequences $(\tau_{1},\tau_{2},\ldots)$. This sum over time sequences gives rise to the additional index $T$, which is absent in Eq. (14). Here, a particular time sequence is generated by
@@
where $\varepsilon_{j+1}$ are random numbers drawn from a uniform distribution in the interval ]0,1]]0,1]] 0 , 1 ]. The amplitudes $A_{j}$, as derived in [72], are now given by
@@
These amplitudes follow from $d_{B_{1},T}(\tau_{j}-0^{+})$, which is evaluated at the bath-contact site $B_{1}$ and just before the time $\tau_{j}$. Therefore, $A_{j}$ only depends on $A_{j-1},\ldots,A_{1}$. If $d_{B_{1}}(\tau_{j}-0^{+})\to 0$, $A_{j}\to 1/2$. Note that $A_{1}=1/2$ by construction. Loosely speaking, $A_{j}$ can be understood as the current “boundary condition” of the density profile, which should not be confused with the actual boundary condition of the Hamiltonian.
While the connection might look quite complicated at first sight, it is rather simple, especially since it expresses the nonequilibrium dynamics in the open system just as a superposition of equilibrium correlation functions in the closed system. As discussed in Ref. [72] in detail, such a connection cannot always hold and requires sufficiently small values of both, $\gamma$ and $\mu$. Thus, one has to check in practise that the result for a given model and a specific choice of the parameters $\gamma$ and $\mu$ does not change when these parameters get smaller and smaller.
A sophisticated criterion for the smallness of the two parameters also exists and has been studied in Ref. [75] in detail. Loosely speaking, the life time of an excitation, which is injected at the boundary of the model, has to be short compared to the time scale $1/\gamma$. More formally, the autocorrelation function at the bath-contact site,
$\langle S_{B_{1}}^{z}(t)S_{B_{1}}^{z}(0)\rangle_{\text{eq}}$ (19)
has to decay on this time scale. For the specific case of periodic boundary conditions, which are used throughout our work, this condition can be satisfied by a sufficiently small $\gamma$ and hence long enough $1/\gamma$. For the other case of open boundary conditions, however, the autocorrelation function does not decay at all for $\Delta>1$, see Ref. [75] and also Refs. [76, 77].
To understand the sophisticated criterion for $\mu$, one has to resort to stochastic unraveling, which underlies the derivation of the connection in Eq. (15) [72]. In this context, $\mu$ has then to be negligible for the deterministic evolution up to the average jump time $1/\gamma$. For the jump probabilities, however, $\mu$ is fully taken into account and gives rise to the amplitues in Eq. (17). Alternatively, by taking a more physical point of view, the smallness of $\mu$ might be seen as a requirement of linear response theory, which then naturally leads to the correlation functions in the connection (15). More details on the validity range and physical meaning can be found in Refs. [72, 75].
@@
In Fig. 2, we illustrate the quality of the connection, by comparing the prediction on the r.h.s. of Eq. (15) to a simulation based on stochastic unraveling for the l.h.s. of Eq. (15). We do so for the model $H^{\prime}$ in Eq. (4) with $\Delta=1.5$, $\Delta^{\prime}=0$, $N=20$, periodic boundary conditions, small coupling $\gamma/J=0.1$, and weak driving $\mu=0.1$. For this set of parameters, the agreement between both sides is remarkable. (Note that stochastic unraveling as such neither requires small $\gamma$ nor weak $\mu$.) While Fig. 2 shows existing data from the literature [72], it already depicts a prediction by the use of classical mechanics, as a main result of our work. This prediction is discussed in the following.
@@ IV Classical limit
with the Kronecker symbol $\delta_{rr^{\prime}}$ and the antisymmetric Levi-Civita tensor $\epsilon_{ijk}$. Here, we write $\hbar$ explicitly, while it is otherwise set to unity. The classical counterpart of our models now results by taking the limit [78]
$S\to\infty\,,\quad\hbar\to 0\,,\quad\hbar\sqrt{S(S+1)}=1\,.$ (21)
@@
$\frac{\text{d}}{\text{d}t}\mathbf{S}_{r}=\frac{\partial H}{\partial\mathbf{S}_{r}}\times\mathbf{S}_{r}\,,$ (23)
$\frac{\partial H}{\partial\mathbf{S}_{r}}=\left(\begin{array}[]{c}S_{r-1}^{x}+S_{r+1}^{x}\\ S_{r-1}^{y}+S_{r+1}^{y}\\ \Delta(S_{r-1}^{z}+S_{r+1}^{z})+\Delta^{\prime}(S_{r-2}^{z}+S_{r+2}^{z})\\ \end{array}\right)\,.$ (24)
Physically, these equations describe the precession of a spin around a magnetic field, which is generated by the interaction with the neighboring spins. Mathematically, they are a coupled set of nonlinear differential equations, which is nonintegrable by means of the Liouville-Arnold theorem, even for $\Delta^{\prime}=0$ [79, 80], which is different to the quantum case $S=1/2$. Therefore, for nontrivial initial conditions, they have to be solved numerically, as we also do here. Throughout our work, we employ a fourth-order Runge-Kutta scheme with a time step $\delta tJ=0.01$, which is small enough to ensure that the total magnetization is well conserved during the time evolution.
@@
where $R_{\text{max}}$ is the number of realizations for different initial conditions, which are randomly drawn from a uniform distribution of points on a unit sphere. Formally, $R_{\text{max}}\to\infty$. In praxis, we average over as many realizations as $R_{\text{max}}={\cal O}(10^{8}-10^{9})$, to ensure that the remaining stochastic fluctuations are low. The need for such an extensive averaging in the numerical simulations is kind of compensated by the fact that the phase space grows only linearly with $N$, in contrast to the exponential growth of the Hilbert space. Due to this fact, we can particularly treat classical systems of quite large size $N$, which lie outside the possibilities of a quantum mechanical treatment. Still, we also consider small $N$ to enable a comparison to quantum mechanics for the same size.
@@
Eventually, coming back to the connection between the dynamics in open and closed systems, Eq. (15), the main idea is to replace the quantum correlation functions by the corresponding classical ones. When we do this strong simplification, the key question of our work is whether or not the time evolution in the open system can be still described accurately. To answer this question in a meaningful way, we first need to rescale the classical time by a factor
@@ V.1 Nearest-neighbor interactions
Using the correlations in Fig. 3, we now calculate the corresponding predictions for the time evolution of the open system, according to Eq. (15). In this calculation, we average over $T_{\text{max}}={\cal O}(10^{4})$ different time sequences, which is enough to obtain sufficiently smooth curves. As already advertised before, Fig. 2 illustrates for $\Delta=1.5$ a convincing agreement with existing simulations [72] on the basis of stochastic unraveling. In Fig. 4, we also depict a comparison for $\Delta=1.0$ and $0.5$. While the agreement for $\Delta=1.0$ is also good, some disagreement is visible for $\Delta=0.5$. Of course, one might expect this disagreement in view of the differences between quantum and classical correlations in Fig. 3 (third row). However, it should be noted that also the quantum prediction differs from the stochastic-unraveling data. This deviation can be traced back to the fact that the ballistic quantum motion for $\Delta=0.5$ partially violates the equilibration assumption, which underlies the derivation of Eq. (15). In other words, while the injected excitation moves away from the boundary, it does not spread homogenoeously over the system [72].
@@ V.3 Larger system sizes
In Fig. 7 (c), we additionally depict a quantity, which we have not discussed so far. This quantity is the injected magnetization by the first bath,
with
as discussed in the appendix in more detail. The injected magnetization is of interest, since from its time derivative, d/$\text{d}/\text{d}t\,\langle\delta S^{z}_{B_{1}}(t)\rangle$, the current in the steady state and then a quantitative value for the diffusion constant can be calculated. As can be seen in Fig. 7 (c), the predictions for $\langle\delta S^{z}_{B_{1}}(t)\rangle$ are rather close to each other. Nevertheless, they still differ. Hence, we obtain two different diffusion coefficients,
$D_{Q}/J\approx 0.99\,,\quad D_{C}/J\approx 0.84\,.$ (29)
Figure 9: Similar data as the one in Fig. 7, but now for a large system size $N=200$. For this $N$, only classical mechanics is feasible. Note that the quality of the linear fit in (b) is much better for $N=200$ (compared to $N=36$), since the relative size of the bath-contact region is substantially smaller.
Because the prediction for the injected magnetization could be more sensitive to short-time deviations, as the ones visible in Fig. 3, we redo the analysis in Fig. 7 for $\gamma/J=0.05$ (instead of $\gamma/J=0.1$). In this way, the relevant time scale is increased. Simultaneously, we use another $\mu=0.05$ (instead of $\mu=0.1$) in Fig. 8. From this data, however, a significant change of the diffusion constants does not result,
$D_{Q}/J\approx 1.01\,,\quad D_{C}/J\approx 0.86\,,$ (30)
which indicates that we do not need to devote special attention to the role of the already small $\gamma$ and $\mu$, to extract quantitative values for transport constants.
It is instructive to compare the values for the diffusion constants in Eq. (30) to other values in the literature. In the classical case, a value of $D_{C}/J\sim 0.6$ has been found for the closed system [47], which is consistent with the value for the open system in Eq. (30). In the quantum case, the value is known with less precision and might be $D_{Q}/J\sim 0.6$ or larger in the closed system [21], while a value of $D_{Q}/J\approx 0.58$ has been found in the open system [23, 24], based on matrix-product-state simulations of the Lindblad equation, yet for a larger $\gamma/J=1$.
@@
$D_{C}/J\approx 0.79\,,$ (31)
which is close to $D_{C}/J\approx 0.84$ for $N=36$, cf. Eq. (29), and indicates a good convergence w.r.t. system size.
@@
Figure 11: Classical data for the system-size dependence of the diffusion constant $D$. While $D$ does not change significantly with $N$ for $\Delta=1.5$, it increases with $N$ for $\Delta=1.0$, which indicates superdiffusion [24]. A power law and a logarithm are depicted as a guide to the eye.
@@
To further substantiate this claim and to conclude on the specific type of anamolous transport, we also depict in Fig. 11 classical data for the system-size dependence of the diffusion constant $D$. At the isotropic point $\Delta=1.0$, $D$ increases with $N$ and indicates superdiffusion [24]. In contrast, at $\Delta=1.5$, $D$ does not change significantly with $N$, as expected for diffusion. While we do not know the specific scaling function for $\Delta=1.0$, we depict in Fig. 11 a power law and a logarithm as a guide to the eye. We should also stress that we cannot exclude a saturation in the large-system limit. For an early discussion of this topic, see Refs. [38, 39, 41, 42].